Find sec θ if tan θ =1/4 and 180° < θ < 270°.
Thank You

Answer by fractalier(6550)

Let q be theta...
We know that tan^2 q + 1 = sec^2 q so that
sec^2 q = 1 + (1/4)^2 = 1 + 1/16 = 17/16
sec q = sqrt(17) / 4
But in the 3rd quadrant sec q, like cos q, is negative, so
sec q = -sqrt(17) / 4